Let   f:  AB be a function. If  f is bijection then   f − 1 is a bijective function from B to A.

Group of answer choices

True

False
Answer #1

solution:

given

  since f : A\rightarrowB be bijective f has invers ^{^{f-1}} :B\rightarrowA

suppose ^{f-1} (b) = ^{f-1} (_{b2}) for some b1 ,b2 in B since f is bijective

f is injective and sub jective since the f is subjective there are elemments

x1 , x2 in a such that b1 = f (x1)

b2 = f (x2)

since f -1 (b1) = f -1 (b2)

we have f -1 (f(x1))= f-1 (f(x2)) holds by defination of invers function

f-1 (f(x1) = x1 and f-1(f(x2)) = x2

\therefore hence x 1 = x2\Rightarrow f(x1) = f(x2) \Rightarrow y1 = y2

y1 = y2\Rightarrow f -1 : B\rightarrowA is injective

note since f is a function ,x1 = x2\Rightarrow f(x1) = f(x2)

since f is a function from A to B for any x\epsilonA there is an y\epsilonB

\therefore y = f(x) \Rightarrow f-1 (y) = f -1(f(x))=x

hence for any x\epsilonA there is an y\epsilonB such that f -1 (y) = x

hence f -1 is a subjection f-1 ia a bijection

the answer is true